# Attractors for the Wave Equation with Nonlinear Energy Critical Sources and Dissipation

### v2

@@ 5.1 Quasi-stability Estimate
$b(t):=2C_{B}C_{1}e^{-\varrho t}e^{C_{2}\int_{0}^{t}d(s,u^{1},u^{2})ds}\quad\mbox{and}\quad c(t):=2C_{2}e^{C_{2}\int_{0}^{t}d(s,u^{1},u^{2})ds}.$ (5.22)
Thus, using that $d(s,u^{1},u^{2})=\|u^{1}_{t}\|^{6}_{6}+\|u^{2}_{t}\|^{6}_{6}\in L^{1}(0,t)$, uniformly in $t>0$ we obtain
$b(t)\in L^{1}(\mathbb{R}^{+})\quad\mbox{and}\quad\lim_{t\rightarrow+\infty}b(t)=0$
and $c(t)$ is bounded on $\mathbb{R}^{+}$ due to $L^{1}(\mathbb{R}^{+})$ integrability of $d(s,u^{1},u^{2})$. The proof of Proposition 5.1 is now complete. ∎
6 Proof of Theorem 2.6: Properties of Attractors
6.1 Proof of Part (I) and Part (III) in Theorem 2.6.
On the virtue of Theorem 2.5, we have established that under the condition $g^{\prime}(0)>0$ the system $(\mathcal{H},S_{t})$ is “quasi-stable”. Therefore, the finite dimension of the attractor $\mathfrak{A}$ (Part (I)) is guaranteed by direct application of Theorem 7.3. As to Part (II) we are in a position to obtain at this point only partial result as stated in (6.2). Indeed, it follows from Proposition 4.1 that the uniform estimate
@@
is valid for any trajectory $\{U(t):t\in\mathbb{R}\}$ lying in the attractor $\mathfrak{A}$. In addition to the quasi-stability property of the system $(\mathcal{H},S_{t})$, using that the function $c(t)$ defined in (5.22) has the property $c_{\infty}=\sup_{t\in\mathbb{R}^{+}}c(t)<\infty$, it follows by direct application of Theorem 7.4 that any complete trajectory $\{U(t):t\in\mathbb{R}\}$ that belongs to the global attractor $\mathfrak{A}$ has the following regularity properties
@@
The proof of Part (III) is an application of Theorem 7.5. First, let $\mathfrak{B}$ be the absorbing set obtained in the Proposition 4.1. From Proposition 5.1, the system $(\mathcal{H},S_{t})$ is quasi-stable on $\mathfrak{B}$. Now, let us consider the spaces $\mathcal{H}_{-s}=H^{-s+1}\times H^{-s}$ with $0<s\leq 1$. Next, we consider the mapping $t\in[0,T]\mapsto S_{t}z\in\mathcal{H}_{-1}=H^{0}\times H^{-1}$. Let $U(t)=S_{t}z=z(t)$ be a weak solution with initial data $z\in\mathfrak{B}$. It follows from Theorem 2.2 that $U_{t}$ has regularity $U_{t}=(u_{t},u_{tt})\in L^{6/5}(0,T;\mathcal{H}_{-1})$. Then, we have
@@
This shows that $t\mapsto S_{t}z$ is Hölder continuous in the extended spaces $\mathcal{H}_{-s}$ for $s\in(0,1]$. Therefore, the existence of a generalized exponential attractor in $\mathcal{H}_{-s}$ with $0<s\leq 1$, follows from Theorem 7.5. ∎
6.2 Maximal Regularity of the Attractor-Completion of the Proof of Part II
To obtain this result we will use arguments similar to those used by Chueshov and Lasiecka in [12, Section 4.2], see aso [29]. The proof follows the following 3 steps. In the first, we prove the $H^{2}$ regularity of the stationary solutions of problem (LABEL:P); in the second, using the regularity of stationary solutions, we establish the smoothness of the trajectories for negative times $t\to$-$\infty$. Finally, in step 3, assuming the regularity obtained in step 2 for an initial data $U(T_{0})\in H^{2}$ with $T_{0}<0$, we propagate this regularity to all $t\in\mathbb{R}.$
@@ 6.2 Maximal Regularity of the Attractor-Completion of the Proof of Part II
$-\Delta u+f(u)=h,\quad u\in H^{1},$ (6.3)
$\|\nabla u\|^{2}+\int_{\Omega}f(u)udx=\int_{\Omega}hudx.$ (6.4)
From Assumption (2.5), we have
$\int_{\Omega}f(u)udx\geq-\frac{\nu}{2}\|u\|^{2}-C_{\nu}|\Omega|\geq-\frac{\nu}{\lambda_{1}}\|\nabla u\|^{2}-C_{\nu}|\Omega|.$ (6.5)
From Hölder inequality, embedding $H^{1}\hookrightarrow H^{0}$, and Young’s inequality, we have
$\left|\int_{\Omega}hudx\right|\leq\|h\|\|u\|\leq\|h\|\frac{1}{\lambda_{1}^{1/2}}\|\nabla u\|\leq\frac{1}{2\omega\lambda_{1}}\|h\|^{2}+\frac{\omega}{2}\|\nabla u\|^{2}.$ (6.6)
Using that $\omega=1-\frac{\nu}{\lambda_{1}}$ and combining (6.4), (6.5), and (6.6), we obtain that
$\|\nabla u\|^{2}\leq\frac{1}{\omega^{2}\lambda_{1}}\|h\|^{2}+\frac{2C_{\nu}|\Omega|}{\omega}.$ (6.7)
Now, multiplying (6.3) by $\Delta u$ and integrating over $\Omega$, we get
@@
From Assumption (2.4) and embedding $H^{2}\hookrightarrow H^{1}$, we get
$-\int_{\Omega}f(u)\Delta udx=\int_{\Omega}f^{\prime}(u)|\nabla u|^{2}dx\geq-K_{f}\|\nabla u\|^{2}.$
where $K_{f}$ arises as in (3.25). From Hölder and Young inequalities, we have
$\left|-\int_{\Omega}h\Delta udx\right|\leq\frac{1}{2\omega}\|h\|^{2}+\frac{\omega}{2}\|\Delta u\|^{2}.$
Substituting the last two inequalities in (6.8) and using (6.7), we conclude that
$||\Delta u||^{2}\leq\frac{1}{\omega^{2}}\|h\|^{2}+\frac{2K_{f}}{\omega}\|\nabla u\|^{2}\leq\frac{1}{\omega^{2}}\|h\|^{2}+\frac{2K_{f}}{\omega}\left(\frac{1}{\omega^{2}\lambda_{1}}\|h\|^{2}+\frac{2C_{\nu}|\Omega|}{\omega}\right)=:\varrho.$ (6.9)
This completes the proof of the lemma. ∎
@@
$\left\{\begin{array}[]{l}w_{tt}-\Delta w+g(w_{t})+f(w)=h,\quad\mbox{in}\quad\Omega\times(s,T)\\ w=0,\quad\mbox{on}\quad\Gamma,\\ (w(s),w_{t}(s))=(u^{*},0),\end{array}\right$. (6.13)
@@
$\left\{\begin{array}[]{l}z_{tt}-\Delta z+g(u_{t})-g(w_{t})=-[\,f(u)-f(w)\,],\quad\mbox{in}\quad\Omega\times(s,T),\\ z=0,\quad\mbox{on}\quad\Gamma,\\ (z(s),z_{t}(s))=(u(s)-u^{*},u_{t}(s)).\end{array}\right$. (6.17)
@@
$\begin{aligned} &\alpha\frac{d}{dt}\int_{\Omega}\nabla w_{t}\nabla wdx+\alpha\|\Delta w(t)\|^{2}+\alpha\int_{\Omega}f^{\prime}(w)|\nabla w|^{2}dx+\alpha\int_{\Omega}h\Delta wdx\\ &=\alpha\|\nabla w_{t}(t)\|^{2}-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx.\end{aligned}$ (6.19)
@@
$\begin{aligned} &\frac{d}{dt}\mathcal{E}_{w}(t)+\alpha\mathcal{E}_{w}(t)+\frac{\alpha}{2}\|\Delta w(t)\|^{2}+\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx\\ &=\frac{3\alpha}{2}\|\nabla w_{t}(t)\|^{2}+\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx+\alpha\int_{\Omega}\nabla w_{t}\nabla wdx.\end{aligned}$ (6.20)
@@
$\begin{aligned} \mathcal{E}_{w}(t)\geq&\;\frac{1}{4}\|\Delta w(t)\|^{2}-\|h\|^{2}-\frac{K_{f}+\alpha^{2}}{2}\|\nabla w\|^{2},\end{aligned}$ (6.21)
@@
$\widetilde{\mathcal{E}}_{w}(t)\geq\frac{1}{4}\|\Delta w(t)\|^{2}.$ (6.22)
@@
$\begin{aligned} &\frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\alpha\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\|\Delta w(t)\|^{2}+\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx\\ &=\alpha\|h\|^{2}+\frac{3\alpha}{2}\|\nabla w_{t}(t)\|^{2}+\alpha\int_{\Omega}\nabla w_{t}\nabla wdx+(K_{f}+\alpha^{2})\int_{\Omega}\nabla w_{t}\nabla wdx\\ &\quad+\frac{\alpha(K_{f}+\alpha^{2})}{2}\|\nabla w(t)\|^{2}+\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx.\end{aligned}$ (6.23)
@@
$\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx$ $\leq$ $C\left[\,1+\|w(t)\|^{3}_{6}\,\right]\|w_{t}(t)\|_{6}\|\nabla w(t)\|^{2}_{6}$
$\leq$ $C\left[\,1+R^{3/2}\,\right]\|w_{t}(t)\|_{6}\|\Delta w(t)\|^{2}$
$\leq$ $\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)+\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t),$
@@
$-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx$ $\leq$ $\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx+\frac{\alpha^{2}}{4}\int_{\Omega}g^{\prime}(w_{t})|\nabla w|^{2}dx$
$\leq$ $\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx+\frac{\alpha^{2}\kappa_{1}}{4}\int_{\Omega}\left[1+|w_{t}|^{4}\right]|\nabla w|^{2}dx$
$\leq$ $\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx+\alpha^{2}C^{\prime}_{R_{1}}\|\Delta w(t)\|^{2}.$
@@
$\frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)+\alpha\left(\frac{1}{2}-\alpha C^{\prime}_{R_{1}}\right)\|\Delta w(t)\|^{2}\leq\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t)+C_{R,R_{1},h},$ (6.24)
@@
$\begin{aligned} \frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)\leq\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t)+C_{R,R_{1},h}.\end{aligned}$ (6.25)
@@
$\begin{aligned} \widetilde{\mathcal{E}}_{w}(t)e^{\frac{\alpha}{2}t}\leq e^{\frac{\alpha}{2}s}\widetilde{\mathcal{E}}_{w}(s)+\int_{s}^{t}\mathbf{d}(\tau,w)e^{\frac{\alpha}{2}\tau}\widetilde{\mathcal{E}}_{w}(\tau)d\tau+\frac{2C_{R_{1},R_{2},h}}{\alpha}e^{\frac{\alpha}{2}t}.\end{aligned}$ (6.26)
@@
$\begin{aligned} \|\Delta w(t)\|^{2}e^{\frac{\alpha}{2}t}\leq 4e^{\|\mathbf{d}\|_{L^{1}}}\left[\,e^{\frac{\alpha}{2}s}\widetilde{\mathcal{E}}_{w}(s)+\frac{2C_{R,R_{1},h}}{\alpha}e^{\frac{\alpha}{2}t}\right].\end{aligned}$ (6.27)
@@
$\widetilde{\mathcal{E}}_{w}(s)\leq C\|\Delta u^{*}\|^{2}+\frac{1}{2}\|h\|^{2}\leq C\varrho+\frac{1}{2}\|h\|^{2}=:\varrho_{0}.$ (6.28)
Thus, taking the limit in (6.27) with $s\to$-$\infty$, we have
$\begin{aligned} \|\Delta w(t)\|^{2}e^{\frac{\alpha}{2}t}\leq\frac{8C_{R,R_{1},h}e^{\|\mathbf{d}\|_{L^{1}(s,t)}}}{\alpha}e^{\frac{\alpha}{2}t}.\end{aligned}$
This implies that
$\begin{aligned} \|\Delta w(t)\|^{2}\leq\frac{8C_{R,R_{1},h}e^{\|\mathbf{d}\|_{L^{1}(s,t)}}}{\alpha}.\end{aligned}$ (6.29)
@@
$\begin{aligned} &E_{z}(t)+\int_{s}^{t}\int_{\Omega}\left[\,g(u_{t})-g(w_{t})\,\right]z_{t}dxd\tau=E_{z}(s)\\ &\quad+\underbrace{K_{f}\int_{s}^{t}\int_{\Omega}z_{t}zdxd\tau}_{\mathrm{J}_{1}}+\underbrace{\frac{1}{2}\int_{s}^{t}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u+(1-\theta)w)[u_{t}+(1-\theta)w_{t}]d\theta|z|^{2}dxd\tau}_{\mathrm{J}_{2}}.\end{aligned}$ (6.30)

### v4

@@ 5.1. Quasi-stability Estimate
(5.21) $b(t):=2C_{B}C_{1}e^{-\varrho t}e^{C_{2}\int_{0}^{t}d(s,u^{1},u^{2})ds}\quad\mbox{and}\quad c(t):=2C_{2}e^{C_{2}\int_{0}^{t}d(s,u^{1},u^{2})ds}.$
Thus, using that $d(s,u^{1},u^{2})=\|u^{1}_{t}\|^{6}_{6}+\|u^{2}_{t}\|^{6}_{6}\in L^{1}(0,t)$, uniformly in $t>0$ we obtain $b(t)\in L^{1}(\mathbb{R}^{+})$, $\lim_{t\rightarrow+\infty}b(t)=0$, and $c(t)$ is bounded on $\mathbb{R}^{+}$ due to $L^{1}(\mathbb{R}^{+})$ integrability of $d(s,u^{1},u^{2})$. The proof of Proposition 5.1 is now complete. ∎
6. Proof of Theorem 2.6: Properties of Attractors
6.1. Proof of Part (I) and Part (III) in Theorem 2.6.
On the virtue of Theorem 2.5, we have established that under the condition $g^{\prime}(0)>0$ the system $(\mathcal{H},S_{t})$ is “quasi-stable”. Therefore, the finite dimension of the attractor $\mathfrak{A}$ (Part (I)) is guaranteed by direct application of [16, Theorem 7.9.6]. As to Part (II) we are in a position to obtain at this point only partial result as stated in (6.2). Indeed, it follows from Proposition 7.1 (Appendix 7) that the uniform estimate
@@
is valid for any trajectory $\{U(t):t\in\mathbb{R}\}$ lying in the attractor $\mathfrak{A}$. In addition to the quasi-stability property of the system $(\mathcal{H},S_{t})$, using that the function $c(t)$ defined in (5.21) has the property $c_{\infty}=\sup_{t\in\mathbb{R}^{+}}c(t)<\infty$, it follows by direct application of [16, Theorem 7.9.8] that any complete trajectory $\{U(t):t\in\mathbb{R}\}$ that belongs to the global attractor $\mathfrak{A}$ has the following regularity properties
@@
The proof of Part (III) is an application of [16, Theorem 7.9.9]. First, let $\mathfrak{B}$ be the absorbing set obtained in the Proposition 7.1. From Proposition 5.1, the system $(\mathcal{H},S_{t})$ is quasi-stable on $\mathfrak{B}$. Now, let us consider the Hilbert spaces $\mathcal{H}_{-s}=H^{-s+1}\times H^{-s}$ with $0<s\leq 1$, where $H^{2s}:=D[(-\Delta)^{s}],$ $s\in\mathbb{R}$, with Dirichlet boundary conditions. Next, we consider the mapping $t\in[0,T]\mapsto S_{t}z\in\mathcal{H}_{-1}=H^{0}\times H^{-1}$. Let $U(t)=S_{t}z=z(t)$ be a weak solution with initial data $z\in\mathfrak{B}$. It follows from Theorem 2.2 that $U_{t}$ has regularity $U_{t}=(u_{t},u_{tt})\in L^{6/5}(0,T;\mathcal{H}_{-1})$. Then, we have
@@
This shows that $t\mapsto S_{t}z$ is Hölder continuous in the extended spaces $\mathcal{H}_{-s}$ for $s\in(0,1]$. Therefore, the existence of a generalized exponential attractor in $\mathcal{H}_{-s}$ with $0<s\leq 1$, follows from [16, Theorem 7.9.9]. ∎
6.2. Maximal Regularity of the Attractor-Completion of the Proof of Part II
To obtain this result we will use arguments similar to those used by Chueshov and Lasiecka in [14, Section 4.2], see aso [32]. The proof follows the following 3 steps. In the first, we prove the $H^{2}$ regularity of the stationary solutions of problem (LABEL:P); in the second, using the regularity of stationary solutions, we establish the smoothness of the trajectories for negative times $t\to$-$\infty$. Finally, in step 3, assuming the regularity obtained in step 2 for an initial data $U(T_{0})\in H^{2}$ with $T_{0}<0$, we propagate this regularity to all $t\in\mathbb{R}.$
@@ 6.2. Maximal Regularity of the Attractor-Completion of the Proof of Part II
Multiplying (4.8) by $\Delta u$ and integrating over $\Omega$, we get
@@
Then, using Assumption (2.4) together with Hölder’s inequality and (4.10), we conclude that
(6.4) $||\Delta u||^{2}\leq\frac{1}{\omega^{2}}\|h\|^{2}+\frac{2K_{f}}{\omega}\|\nabla u\|^{2}\leq\frac{1}{\omega^{2}}\|h\|^{2}+\frac{2K_{f}}{\omega}\left(\frac{1}{\omega^{2}\lambda_{1}}\|h\|^{2}+\frac{2C_{\nu}|\Omega|}{\omega}\right)=:\varrho,$
where $K_{f}$ arises as in (3.18). This completes the proof of the lemma. ∎
@@
(6.8) $\left\{\begin{array}[]{l}w_{tt}-\Delta w+g(w_{t})+f(w)=h,\quad\mbox{in}\quad\Omega\times(s,T)\\ w=0,\quad\mbox{on}\quad\Gamma,\\ (w(s),w_{t}(s))=(u^{*},0),\end{array}\right$.
@@
(6.12) $\left\{\begin{array}[]{l}z_{tt}-\Delta z+g(u_{t})-g(w_{t})=-[\,f(u)-f(w)\,],\quad\mbox{in}\quad\Omega\times(s,T),\\ z=0,\quad\mbox{on}\quad\Gamma,\\ (z(s),z_{t}(s))=(u(s)-u^{*},u_{t}(s)).\end{array}\right$.
@@
(6.14) $\begin{aligned} &\alpha\frac{d}{dt}\int_{\Omega}\nabla w_{t}\nabla wdx+\alpha\|\Delta w(t)\|^{2}+\alpha\int_{\Omega}f^{\prime}(w)|\nabla w|^{2}dx+\alpha\int_{\Omega}h\Delta wdx\\ &=\alpha\|\nabla w_{t}(t)\|^{2}-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx.\end{aligned}$
@@
(6.15) $\begin{aligned} &\frac{d}{dt}\mathcal{E}_{w}(t)+\alpha\mathcal{E}_{w}(t)+\frac{\alpha}{2}\|\Delta w(t)\|^{2}+\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx\\ &=\frac{3\alpha}{2}\|\nabla w_{t}(t)\|^{2}+\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx+\alpha\int_{\Omega}\nabla w_{t}\nabla wdx.\end{aligned}$
@@
(6.16) $\begin{aligned} \mathcal{E}_{w}(t)\geq&\;\frac{1}{4}\|\Delta w(t)\|^{2}-\|h\|^{2}-\frac{K_{f}+\alpha^{2}}{2}\|\nabla w\|^{2},\end{aligned}$
@@
(6.17) $\widetilde{\mathcal{E}}_{w}(t)\geq\frac{1}{4}\|\Delta w(t)\|^{2}.$
@@
(6.18) $\begin{aligned} &\frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\alpha\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\|\Delta w(t)\|^{2}+\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx\\ &=\alpha\|h\|^{2}+\frac{3\alpha}{2}\|\nabla w_{t}(t)\|^{2}+\alpha\int_{\Omega}\nabla w_{t}\nabla wdx+(K_{f}+\alpha^{2})\int_{\Omega}\nabla w_{t}\nabla wdx\\ &\quad+\frac{\alpha(K_{f}+\alpha^{2})}{2}\|\nabla w(t)\|^{2}+\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx.\end{aligned}$
@@
$\frac{1}{2}\int_{\Omega}f^{\prime\prime}(w)w_{t}|\nabla w|^{2}dx$ $\leq$ $\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)+\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t),$
@@
$-\alpha\int_{\Omega}g^{\prime}(w_{t})\nabla w\nabla w_{t}dx$ $\leq$ $\int_{\Omega}g^{\prime}(w_{t})|\nabla w_{t}|^{2}dx+\alpha^{2}C^{\prime}_{R_{1}}\|\Delta w(t)\|^{2}.$
@@
(6.19) $\frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)+\alpha\left(\frac{1}{2}-\alpha C^{\prime}_{R_{1}}\right)\|\Delta w(t)\|^{2}\leq\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t)+C_{R,R_{1},h},$
@@
(6.20) $\begin{aligned} \frac{d}{dt}\widetilde{\mathcal{E}}_{w}(t)+\frac{\alpha}{2}\widetilde{\mathcal{E}}_{w}(t)\leq\mathbf{d}(t,w)\widetilde{\mathcal{E}}_{w}(t)+C_{R,R_{1},h}.\end{aligned}$
@@
(6.21) $\begin{aligned} \widetilde{\mathcal{E}}_{w}(t)e^{\frac{\alpha}{2}t}\leq e^{\frac{\alpha}{2}s}\widetilde{\mathcal{E}}_{w}(s)+\int_{s}^{t}\mathbf{d}(\tau,w)e^{\frac{\alpha}{2}\tau}\widetilde{\mathcal{E}}_{w}(\tau)d\tau+\frac{2C_{R_{1},R_{2},h}}{\alpha}e^{\frac{\alpha}{2}t}.\end{aligned}$
@@
(6.22) $\begin{aligned} \|\Delta w(t)\|^{2}e^{\frac{\alpha}{2}t}\leq 4e^{\|\mathbf{d}\|_{L^{1}}}\left[\,e^{\frac{\alpha}{2}s}\widetilde{\mathcal{E}}_{w}(s)+\frac{2C_{R,R_{1},h}}{\alpha}e^{\frac{\alpha}{2}t}\right].\end{aligned}$
@@
(6.23) $\widetilde{\mathcal{E}}_{w}(s)\leq C\|\Delta u^{*}\|^{2}+\frac{1}{2}\|h\|^{2}\leq C\varrho+\frac{1}{2}\|h\|^{2}=:\varrho_{0}.$
Thus, taking the limit in (6.22) with $s\to$-$\infty$, we obtain
(6.24) $\begin{aligned} \|\Delta w(t)\|^{2}\leq\frac{8C_{R,R_{1},h}e^{\|\mathbf{d}\|_{L^{1}(s,t)}}}{\alpha}.\end{aligned}$
@@
(6.25) $\begin{aligned} &E_{z}(t)+\int_{s}^{t}\int_{\Omega}\left[\,g(u_{t})-g(w_{t})\,\right]z_{t}dxd\tau=E_{z}(s)\\ &\quad+\underbrace{K_{f}\int_{s}^{t}\int_{\Omega}z_{t}zdxd\tau}_{\mathrm{J}_{1}}+\underbrace{\frac{1}{2}\int_{s}^{t}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u+(1-\theta)w)[u_{t}+(1-\theta)w_{t}]d\theta|z|^{2}dxd\tau}_{\mathrm{J}_{2}}.\end{aligned}$